# AED-PADA:Improving Generalizability of Adversarial Example Detection via Principal Adversarial Domain Adaptation

### v1

@@ 3.3 Adversarial Domain Clustering
After the acquisition of ADs, we obtain $\Gamma_{i}=Enc(D_{i})$ as AD of $Adv_{i}$. For $\mathbb{O}$, we construct a set of ADs, =$\mathbb{R}=\{\Gamma_{1},\Gamma_{2},\ldots,\Gamma_{M}\}$. Since different ADs tend to distribute differently in the feature space and many of them possess various portions of overlaps, it is quite difficult to directly select the most representative ADs from scratch. Then, it is vital to explore the similarities among different ADs.
@@
To address this issue, we intend to perform the selection via two steps, i.e., clustering to assess the similarities of ADs and selecting the most representative ADs from each cluster. Then, we construct a viable solution named Adversarial Domain Clustering, as shown in Fig. 2(a). This strategy groups ADs into different clusters, by ensuring that the similarity among ADs within the same cluster is maximized, while the similarity among ADs across different clusters is minimized.
@@
Since the samples to be clustered here are collections of features, rather than individual data points, traditional clustering methods such as K-Means [?] cannot be directly applied. Since spectral clustering [?] only requires the similarity matrix among samples, it is utilized to construct the clustering step in our AD clustering.
@@
For the estimation of the similarity matrix $W\in R^{M\times M}$ in spectral clustering, which represents the similarities between different ADs, we propose Adversarial Domain Similarity Measurement (ADSM) based on Jensen-Shannon divergence (JSD) [?], which quantifies the similarity between two probability distributions. To compute the similarities, we transform each $\Gamma_{i}$ in $\mathbb{R}$ to $\overline{\Gamma_{i}}$ by converting $\Gamma_{i}$ into probabilities and performing normalizations. Since a smaller value of JSD between two ADs implies a smaller discrepancy in their probability distributions, $\text{ADSM}(\Gamma_{i},\Gamma_{j})$ can be computed via
$\text{ADSM}(\Gamma_{i},\Gamma_{j})=\frac{1}{\text{JSD}(\overline{\Gamma_{i}},\overline{\Gamma_{j}})}.$ (3)
@@
By letting the element $W_{i,j}$ at the $i$-th row and $j$-th column refer to the similarity between $\Gamma_{i}$ and $\Gamma_{j}$, $W_{i,j}$ can be calculated by
$W_{i,j}=\left\{\begin{array}[]{cc}\text{ADSM}(\Gamma_{i},\Gamma_{j})&i\neq j\\ 0&i=j\end{array}.\right.$ (4)
@@
Since the spectral clustering cannot automatically determine the optimal number of clusters, to evaluate the clustering performance and estimate the optimal number of clusters, the Calinski-Harabasz score (CH score) [?], which requires no knowledge of the cluster shape, is utilized. Note that it measures both the within-cluster and between-cluster distances, thereby offering a more comprehensive view of the clustering performance. With the help of the CH score, our AD clustering can automatically group the ADs into optimal number of clusters.
3.4 Principal Adversarial Domains Selection
@@
After similar ADs are clustered, then the selection step can be performed. To form the most effective Principal Adversarial Domains (PADs), it is vital to select proper ADs from different clusters. Thus, we propose the Coverage of Entire Feature Space metric (CEFS) to guide the PADs selection process.
@@
CEFS is a ratio-based metric which contains two aspects, Intra-Domain Dispersion (IDD) and Discrepancy between Adversarial Domains (DAD). IDD represents the dispersion among features within each AD, where a higher value indicates a larger coverage of the feature space. For any AD, $\Gamma_{I}=\{{\Gamma_{I}}_{1},{\Gamma_{I}}_{2},\ldots,{\Gamma_{I}}_{v}\}\in R^{v\times d}$, where $v$ and $d$ denote the number and dimension of features, respectively, IDD can be computed by
$\text{IDD}(x)=\frac{1}{v}\sum_{i=1}^{v}\left\|\frac{x_{i}}{\|x_{i}\|_{2}}-\frac{1}{v}\sum_{j=1}^{v}\frac{x_{j}}{\|x_{j}\|_{2}}\right\|_{2}.$ (5)
@@
DAD represents the discrepancies between the two selected ADs, this discrepancy can be quantified using a distance metric $Dist$, such as Kullback-Leibler divergence [?] or Maximum Mean Discrepancy (MMD) [?]. A lower DAD value indicates a larger similarity of the two selected ADs. DAD can be calculated as,
$\text{DAD}(\Gamma_{i},\Gamma_{j})=Dist(\overline{\Gamma_{i}},\overline{\Gamma_{j}}).$ (6)
@@
$\text{CEFS}=\frac{\sum_{i}^{K}\text{IDD}(\Gamma_{i})}{\text{DAD}((\Gamma_{1}\parallel\cdots\parallel\Gamma_{K}),(\Gamma_{1}\parallel\cdots\parallel\Gamma_{M}))},$ (7)
where $K$ and $M$ is the number of ADs in PADs and the number of ADs in the AD set $\mathbb{R}$ respectively, and $\parallel$ is a concatenation operation. A higher CEFS value represents a larger coverage of the entire feature space. We utilize CEFS to select PADs, which can give a larger coverage of the feature space with the same number of ADs. PADs actually enhance the probability of capturing the location of unseen ADs, thereby improving the generalization performance.
3.5 Principal Adversarial Domain Adaptation
@@
Ideally, the inputs of PADA comprise source data and unseen target data. The source data contains the adversarial examples from PADs. Since the unseen target data is unavailable, we can only utilize $\mathcal{O}$, the entire set of adversarial examples, as a proxy.
PADA consists of 3 sequential components, feature extraction, feature alignment and classification. PADA is compatible with various widely used multi-source domain adaptation (MDA) methods [???].
@@
Unfortunately, existing MDA methods only extract spatial features. [??] indicate that adversarial perturbations are more likely to be concealed in the high-frequency information of images. To capture more comprehensive features of adversarial examples, we propose an adversarial feature enhancement (AFE) module to be the feature extraction component. AFE contains both the spatial feature extraction $\mathcal{S}$ and frequency feature extraction $\mathcal{F}$ branches. For frequency feature extraction, we design a plug-and-play high-pass filters (HPF) operation to enhance the high-frequency signals and remove low-frequency components of inputs. More details of HPF can be found in the supplementary materials. Both $\mathcal{S}$ and $\mathcal{F}$ employ $Enc$ from Sec. 3.2. Then, the enhanced adversarial features, denoted as $r(x)=[\mathcal{S}(x),\mathcal{F}(x)]$, are fed into the future module.
4 Experiments
4.1 Experimental Setups
DNN backbones and datasets.
@@ 4.1 Experimental Setups
We evaluate the performance of the proposed method, by employing two widely used DNN architectures, i.e., ResNet-18  [?] and VGG-16 [?], on two popular datasets CIFAR-10 [?] and SVHN [?], by following [?].
Baseline attack methods.

### v2

@@ 3.3. Adversarial Domain Clustering
After the acquisition of ADs, we obtain $\mathbb{H}^{m}=Enc(\mathbb{D}^{m})$ as AD of the corresponding adversarial attack $\psi^{m}$. For $\mathcal{D}$, we construct a set of ADs, =$\mathcal{H}=\{\mathbb{H}^{m}\}_{m=1}^{M}$. Since different ADs tend to distribute differently in the feature space and many of them possess various portions of overlaps, it is quite difficult to directly select the most representative ADs from scratch. Then, it is vital to explore the similarities among different ADs.
@@
To address this issue, we intend to perform the selection via two steps, i.e., clustering to assess the similarities of ADs and selecting the most representative ADs from each cluster. Then, we construct a viable solution named Adversarial Domain Clustering, as shown in Fig. 2(a). This strategy groups ADs into different clusters, by ensuring that the similarity among ADs within the same cluster is maximized, while the similarity among ADs across different clusters is minimized. AD clustering avoids the redundancy and additional costs by preventing the repeated selection of similar attack methods, and it facilitates the selection of the most representative ADs in the subsequent steps.
@@
Since the samples to be clustered here are collections of features, rather than individual data points, traditional clustering methods such as K-Means (kmeans, ) cannot be directly applied. Since spectral clustering (NgJW01, ) only requires the similarity matrix among samples, it is utilized to construct the clustering step in our AD clustering.
@@
For the estimation of the similarity matrix $W\in R^{M\times M}$ in spectral clustering, which represents the similarities between different ADs, we propose Adversarial Domain Similarity Measurement (ADSM) based on Jensen-Shannon divergence (JSD) (ErvenH14, ), which quantifies the similarity between two probability distributions. To compute the similarities, we transform each $\mathbb{H}^{i}$ in $\mathcal{H}$ to $\overline{\mathbb{H}^{i}}$ by converting $\mathbb{H}^{i}$ into probabilities and performing normalizations. Since a smaller value of JSD between two ADs implies a smaller discrepancy in their probability distributions, $\text{ADSM}(\mathbb{H}^{i},\mathbb{H}^{j})$ can be computed via
(3) $\text{ADSM}(\mathbb{H}^{i},\mathbb{H}^{j})=\frac{1}{\text{JSD}(\overline{\mathbb{H}^{i}},\overline{\mathbb{H}^{j}})}.$
@@
By letting the element $W_{i,j}$ at the $i$-th row and $j$-th column refer to the similarity between $\mathbb{H}^{i}$ and $\mathbb{H}^{j}$, $W_{i,j}$ can be calculated by
(4) $W_{i,j}=\left\{\begin{array}[]{cc}\text{ADSM}(\mathbb{H}^{i},\mathbb{H}^{j})&i\neq j\\ 0&i=j\end{array}.\right.$
@@
Since the spectral clustering cannot automatically determine the optimal number of clusters, the Calinski-Harabasz score (CH score) (mann2023proposed, ), which requires no knowledge of the cluster shape, is utilized to evaluate the clustering performance and estimate the optimal number of clusters. Note that it measures both the within-cluster and between-cluster distances, thereby offering a more comprehensive view of the clustering performance. CH score is calculated by
(5) $\text{CH}(K)=\frac{Tr(BC_{K})/(K-1)}{Tr(WC_{K})/(N-K)},$
where $K$ and $N$ is the number of clusters and data respectively, $BC_{K}$ denotes between-cluster covariance matrix, $WC_{K}$ denotes within-cluster covariance matrix, and $Tr(\cdot)$ denotes the trace of the matrix.
Higher value of $\text{CH}(K)$ indicates better clustering. We posit that the inherent structure of the entire feature space composed of different ADs is highly complex. Given that the CH score often awards the highest evaluation when cluster numbers $K=2$, we choose to commence our consideration from the scenario where $K=3$ in this paper. With the help of the CH score, our AD clustering can automatically group the ADs into optimal number of clusters.
3.4. Principal Adversarial Domains Selection
@@
After similar ADs are clustered, then the selection step can be performed. To form the most effective Principal Adversarial Domains (PADs), it is vital to select appropriate ADs from different clusters. Thus, we propose the Coverage of Entire Feature Space metric (CEFS) to guide the PADs selection process.
@@
CEFS is a ratio-based metric which contains two aspects, Intra-Domain Dispersion (IDD) and Discrepancy between Adversarial Domains (DAD). IDD represents the dispersion among features within each AD, where a higher value indicates a larger coverage of the feature space. For any AD, $\mathbb{H}^{i}=\{{h_{1}^{i}},{h_{2}^{i}},\ldots,{h_{v}^{i}}\}\in R^{v\times d}$, where $v$ and $d$ denote the number and dimension of features in $\mathbb{H}^{i}$, respectively, IDD can be computed by
(6) $\text{IDD}(x)=\frac{1}{v}\sum_{i=1}^{v}\left\|\frac{x_{i}}{\|x_{i}\|_{2}}-\frac{1}{v}\sum_{j=1}^{v}\frac{x_{j}}{\|x_{j}\|_{2}}\right\|_{2}.$
@@
DAD represents the discrepancies between the two selected ADs, this discrepancy can be quantified using a distance metric $Dist(\cdot,\cdot)$, such as Kullback-Leibler divergence (kl-divergence, ) or Maximum Mean Discrepancy (MMD) (mmd, ). A lower DAD value indicates a greater similarity between the two selected ADs. DAD can be calculated as,
(7) $\text{DAD}(\mathbb{H}^{i},\mathbb{H}^{j})=Dist(\overline{\mathbb{H}^{i}},\overline{\mathbb{H}^{j}}).$
@@
(8) $\text{CEFS}=\frac{\sum_{i=1}^{K}\text{IDD}(\mathbb{H}^{i})}{\text{DAD}((\mathbb{H}^{1}\parallel\cdots\parallel\mathbb{H}^{K}),(\mathbb{H}^{1}\parallel\cdots\parallel\mathbb{H}^{M}))},$
where $K$ and $M$ is the number of ADs in PADs and the number of ADs in the AD set $\mathcal{H}$ respectively, and $\parallel$ denotes a concatenation operation.
CEFS is a ratio-based metric to quantify the coverage of selected ADs within the entire feature space (EFS). In Eq. (8), the numerator represents Intra-Domain Dispersion (IDD), indicates feature dispersion within each AD. The denominator measures the discrepancy between the selected ADs and EFS. As CEFS increases, the numerator grows, indicating a larger feature space for each AD, and the denominator decreases, suggesting a greater similarity between the selected ADs and the EFS. Consequently, the selected ADs have a larger coverage of the EFS, increasing the likelihood of capturing unseen ADs. We utilize CEFS to select PADs, which can give a larger coverage of the feature space with the same number of ADs. PADs actually enhance the probability of capturing the location of unseen ADs, thereby improving the generalization performance.
3.5. Training process of Principal Adversarial Domains Identification
The process of Principal Adversarial Domains Identification (PADI) consists of Adversarial Domain Acquisition (AD Acquisition), Adversarial Domain Clustering (AD Clustering) and Principal Adversarial Domains Selection (PADs Selection). The training process of Principal Adversarial Domains Identification is described in Algorithm 1.
Algorithm 1 Training process of Principal Adversarial Domains Identification
1:$\mathcal{D}_{acq},\mathcal{D}_{clu}$.
2:Principal Adversarial Domains (PADs).
3$\triangleright$ Adversarial Domain Acquisition
4:for $x_{i}^{m_{i}}$ $\mathbf{in}$ $\mathcal{D}_{acq}$ do
5:     Augment $x_{i}^{m_{i}}$ to be $(\tilde{x}_{2i-1},\tilde{x}_{2i})$ with augmentation $Norm$
6:     $z_{i}=Enc(Proj(\tilde{x}_{i}))$
7:     Get $\mathbf{Enc}$, $\mathbf{Proj}$ via minimizing $\mathcal{L}$ in Eq. (1)
8:end for
9$\triangleright$ Adversarial Domain Clustering
10:Generate the AD pool $\mathcal{H}_{clu}$ based on $\mathcal{D}_{clu}$
11:Generate the similarity matrix $W$ by Eq. (3)
12:for  $k$ $\mathbf{in}$ [3,$M$do
13:     $r_{k}=SC(W,k)$triangleright $r_{k}$ is the clustering result via Spectral clustering.
14:     Get $\text{CH}(k)$ by Eq. (5)
15:end for
16:$K=\operatorname*{arg\,max}_{k}(\text{CH}(k))$triangleright $K$ is the best number of clusters.
17:$r_{K}=SC(W,K)=\{clu_{i}\}_{i=1}^{K}$triangleright $r_{K}$ is the best clustering result which contains $K$ clusters.
18$\triangleright$ Principal Adversarial Domains Selection
19:$C=\Pi_{i=1}^{K}n_{i}$triangleright $n_{i}$ is the number of adversarial attacks in $clu_{i}$, and $\Sigma_{i=1}^{K}n_{i}=M.$
20:Generate combination list $L=\{l_{i}\}_{i=1}^{C}$triangleright $L$ is the list of all combinations in $r_{K}$, each $L_{i}$ has $K$ types of attacks.
21:for $l_{i}$ $\mathbf{in}$ $L$  do
22:     Generate $\mathcal{H}_{i}=\{\mathbb{H}_{i}^{j}\}_{j=1}^{K}$ based on $l_{i}$$\triangleright$ $\mathcal{H}_{i}$ is the list of ADs based on $l_{i}$
23:     Get CEFS($\mathcal{H}_{i}$) by Eq. (8)
24:end for
25:$\text{PADs}=\operatorname*{arg\,max}_{l_{i}}(\text{CEFS}(\mathcal{H}_{i}))$
For the original dataset $\mathbb{D}$, we divide it into two non-overlapping datasets $\mathbb{D}_{acq}$ and $\mathbb{D}_{clu}$, which are used for AD Acquisition and AD Clustering, respectively. The adversarial examples sets for AD Acquisition are denoted as $\mathcal{D}_{acq}$. $\mathcal{D}_{acq}=\{\mathbb{D}_{acq}^{m}\}_{m=1}^{M},\mathbb{D}_{acq}^{m}=\{(x_{i}^{m},y_{i}^{m})\}_{i=1}^{N_{m}}$, where $M$ is the number of adversarial attacks, $\mathbb{D}_{acq}^{m}$ is the adversarial examples set generated by the $m$-th adversarial attack, and $N_{m}$ is the cardinality of $\mathbb{D}_{acq}^{m}$.
Likewise, The adversarial examples sets for AD Clustering are denoted as $\mathcal{D}_{clu}$. $\mathcal{D}_{clu}=\{\mathbb{D}_{clu}^{q}\}_{q=1}^{M}$, $\mathbb{D}_{clu}^{q}=\{(x_{i}^{q},y_{i}^{q})\}_{i=1}^{N_{q}}$, where $M$ is the number of adversarial attacks, $\mathbb{D}_{clu}^{q}$ is the adversarial examples set generated by the $q$-th adversarial attack, and $N_{q}$ is the cardinality of $\mathbb{D}_{clu}^{q}$. The rationale behind this strategy is to prevent the model after contrastive learning from overfitting to the data on AD acquisition, which leads to subpar performance in AD Clustering and PAD Confirming.
3.6. Principal Adversarial Domain Adaptation
@@
Ideally, the inputs of PADA comprise source data and unseen target data. The source data, denoted as, $Src=\{Src_{i}\}_{i=1}^{K}$, consists of an equal number of benign examples and adversarial examples, where the adversarial examples contain $K$ types of adversarial examples determined by PADs. Since the unseen target data is unavailable, we can only use the training data as a proxy. The proxy data, denoted as $PD$, also contains an equal number of benign examples and adversarial examples, drawn from the training benign set $\mathbb{D}$ and the training adversarial set $\mathcal{D}$, respectively. $\mathcal{D}$ contains $M$ types of adversarial examples. During the PADI stage, we select the most representative ADs, i.e., $K$ types of ADs from these $M$ types to form the PADs. Consequently, employing $\mathcal{D}$ as a proxy for unseen target data serves two key purposes. Firstly, it prevents overlap between the source data used for training and the unseen target data used for testing. Secondly, the PADA process compels the transfer of specific knowledge from PADs to the more extensive set $\mathcal{D}$. This transfer aims to boost the generalization capabilities of networks of PADA, with the goal of enhancing their performance when testing the unseen target data.
PADA consists of three sequential components, feature extraction, feature alignment and classification. The framework of PADA is compatible with various widely used Multi-source Unsupervised Domain Adaptation (MUDA) methods (moment-mda, ; DARN, ; mdmn, ; aaai-mda, ). The experimental results indicate that our PADA possesses excellent generalization capabilities based on various existing MUDA methods. Due to the simplicity and effectiveness of MFSAN (aaai-mda, ), along with its superior detection performance compared to other MUDA methods, we select MFSAN as the basic MUDA method for our PADA.
Figure 3. The three kernels of PEF employed to capture the perturbation signals hidden in the adversarial inputs.
@@
In the feature extraction component, unfortunately, existing MUDA methods including MFSAN, only extract spatial features. (GeirhosRMBWB19, ; WangWHX20, ; FanLCZG21, ) indicate that the high-frequency component of an image plays a crucial role in the prediction of deep neural network. Adversarial perturbations are more likely to be concealed in the high-frequency information of images. To capture more comprehensive features of adversarial examples, we propose an adversarial feature enhancement (AFE) module as the feature extraction component. AFE contains both the spatial feature extraction $\mathcal{S}$ and frequency feature extraction $\mathcal{F}$ branches. For frequency feature extraction, we design perturbations extraction filters (PEF), which is a plug-and-play operation based on Spatial Rich Model (SRM) (SRM, ) to capture subtle perturbation signals hidden in the adversarial examples.
SRM typically uses 30 basic kernels to capture textures and discontinuities of images, and has been employed in image forensics to detect subtle and irregular manipulation or hidden information. Subsequent research (SRM_detection, ) indicates that in image manipulation detection, employing only three kernels can achieve considerable detection performance, and more many basic kernels does not further enhance performance. Essentially, these kernels are high-pass filters, which enhance the high-frequency signals and remove the low-frequency components of the inputs. Adversarial perturbations are typically hidden within the high-frequency information of an image. Therefore, our PEF uses the same three kernels to extract subtle perturbation signals. As shown in Fig. 3, PEF consists of three filters which are implemented by convolution kernels with fixed parameters. We set the kernel size of PEF to be $5\times 5\times 3$, and the output channel size of PEF is 3. Both $\mathcal{S}$ and $\mathcal{F}$ employ $Enc$ from Sec. 3.2, aligned with the threat models, specifically ResNet-18 or VGG-16. Then, the enhanced adversarial features, denoted as $r(x)=[\mathcal{S}(x),\mathcal{F}(x)]$, are fed into the next module.
In the alignment component, we assign the specific network $Q_{i}$ to map each source domain $Src_{i}$ and proxy domain $PD$ to the different feature space, and utilizes MMD as the distance metric to align them. $L_{d}$ is used to align the features between source domain and proxy domain,
(9) $L_{d}=\frac{1}{K}\sum_{i=1}^{K}\text{MMD}(Q_{i}(r(Src_{i})),Q_{i}(r(PD))),$
where $r$ is the enhanced adversarial features after AFE, $K$ is the number of source domains. each source domain is associated with the corresponding network $Q_{i}$.
In the classification component, we utilize the cross-entropy loss $L_{cls}$ to ensure the correct classification. Given that different domain-specific classifiers $C_{i}$ are trained on their respective source domains, resulting in significant discrepancies among their predictions for the same proxy domain. To address this, $L_{disc}$ is used to minimize the differences among the predictions of various classifiers.
(10) $L_{disc}=$ $\frac{2}{K\times(K-1)}\sum_{j=1}^{K-1}\sum_{i=j+1}^{K}\mathbb{E}_{x\sim PD}|C_{i}(Q_{i}(r(x)))-C_{j}(Q_{j}(r(x)))|.$
Overall, the total loss is formulated as follow, $\lambda$ and $\gamma$ are hyperparameters used to adjust the weights of $L_{d}$ and $L_{disc}$, respectively.
(11) $L_{total}=L_{cls}+\lambda L_{d}+\gamma L_{disc}.$
4. Experiments
4.1. Experimental Setups
4.1.1. DNN backbones
@@
We evaluate the performance of the proposed method, by employing two widely used DNN architectures, i.e., ResNet-18  (HeZRS16, ) and VGG-16 (SimonyanZ14a, ), according to (SID, ).
4.1.2. Datasets
We evaluate the performance of the proposed method on three popular datasets, CIFAR-10 (cifar10, ), SVHN (SVHN, ) and ImageNet (imagenet, ). All the images in ImageNet are resized to $224\times 224\times 3$ via pre-processing.
As shown in Table 1, each of the three datasets is divided into three category-balanced and non-overlapping subsets: Train-acq-src, Train-clu-pro and Test. The training data for AD Acquisition in the PADI stage and the source data in the PADA stage are selected from Train-acq-src. Similarly, the training data for AD Clustering in the PADI stage and the proxy data in the PADA stage are selected from Train-clu-pro.
For CIFAR-10, we divide the official CIFAR-10 training set into two halves, each of which contains 25,000 images, to form Train-acq-src and Train-clu-pro. Besides, we form Test with all the 10,000 images in the official CIFAR-10 testing set. For SVHN, we set the number of Train-acq-src and Train-clu-pro to 20,000 instead of 25,000, due to the presence of category imbalance in the official SVHN training dataset, to construct a category-balanced training dataset. Test of SVHN also consists of 10,000 images, which are randomly selected from the official SVHN testing set. For ImageNet, We divide the official ImageNet (ILSVRC2012) validation set, which consists of 50,000 images, into two parts, 40,000 images for training and 10,000 images for Test. The 40,000 training images are then equally divided into two subsets, Train-acq-src and Train-clu-pro, each of which contains 20,000 images.
Ten types of training adversarial examples are created based on the benign examples from the Train-acq-src and Train-clu-pro subsets. Subsequently, another seven attacks are applied to generate the testing adversarial examples on the Test set. These specific adversarial attack methods are detailed in Sec. 4.1.4. Note that the attack methods for training and testing are entirely different, ensuring that the training data and testing data are mutually exclusive.
4.1.3. Data splitting strategy for the training stage
Table 2 presents the data splitting strategy for the training stage of our AED-PADA. AED-PADA contains two stages during training. In the PADI stage, to improve the efficiency, we randomly select 10,000 images from each type of adversarial attacks and their corresponding benign images from the Train-acq-src for supervised contrastive learning. To alleviate the overfitting problem, we select 10,000 images per attack from Train-clu-pro for AD Clustering and PAD Confirming.
In the PADA stage, the inputs of PADA are the images from the source domains and the unseen target domains. Each source domain contains 10,000 samples, with an equal split of 5,000 benign examples and 5,000 adversarial examples, which are all randomly selected from Train-acq-src. Since the data from unseen domains is unavailable, we can only utilize the training examples as a proxy, which are generated from all the 10 types of training attacks. Specifically, the proxy domain contains 10,000 samples, i.e., 5,000 benign examples and 5,000 adversarial examples, which are all randomly selected from the Train-clu-pro. Note that the adversarial examples in the proxy domain are obtained via 10 different training attacks, with each attack providing 500 examples. This specific arrangement has two benefits. Firstly, it ensures zero overlap between the source and proxy domains during training, to avoid data leakage. Secondly, it enhances the diversity of the training data, to further benefit the generalization of the detection model.
Table 1. Datasets for training and testing.
Dataset Train-acq-src Train-clu-pro Test
benign adv benign adv benign adv
CIFAR-10 $25,000$ $10\times 25,000$ $25,000$ $10\times 25,000$ $10,000$ $7\times 10,000$
SVHN $20,000$ $10\times 20,000$ $20,000$ $10\times 20,000$ $10,000$ $7\times 10,000$
ImageNet $20,000$ $10\times 20,000$ $20,000$ $10\times 20,000$ $10,000$ $7\times 10,000$
Table 2. Data splitting for the training stage of our AED-PADA.
Training stage benign examples adversarial examples
PADI AD Acquisition - $10,000$/attack
AD Clustering - $10,000$/attack
PADA source domain $5,000$/source $5,000$/source
proxy domain $5,000$ $5,000$
4.1.4. Baseline adversarial attack methods